# Collapsing in polygonal dynamics

### v2

@@ 5 Two examples of dynamics in $\mathbb{P}^{1}$
In this section, we study two polygonal dynamics previously defined in other papers. For any of them, we start by deriving their scaling symmetries, and then apply theorem 4.2 to get their infinitesimal monodromies. From this, we get what are their possible collapse points.
@@ 5.1 The leapfrog map
The leapfrog map is defined in [11, §5.2] as an analogue of the pentagram map in $\mathbb{P}^{1}$, and the authors notice that it seems to be the hyperbolic counterpart of the polygon recutting defined in [3] (which would be the euclidean case). For a matter of internal coherence, the notations of [11] are changed to fit the ones from here.
@@
with the same monodromy, which we can intertwine this way
@@
There is a scaling symmetry acting by:
@@
We will express the system of coordinates $(o^{-},o)$ with the help of two other systems, and then get the result. According to [11, Prop. 5.12], we have the coordinate systems $(x,y)$ and $(p,r)$, defined by:
@@
They are linked by the following relations [^11] The original paper forgot the minus signs, but it doesn’t impact their reasoning.:
@@
But according to [11, Rmk. 3.5], there is a scaling symmetry (actually working for all of their generalised pentagram maps) defined for any $t\in\mathbb{G}_{m}$ by
@@
$M^{\prime}=\sum_{i=1}^{n}\frac{1}{s_{i}^{-}-s_{i}}\begin{pmatrix}s_{i}^{-}&-s_{i}^{-}s_{i}\\ 1&-s_{i}\end{pmatrix}$,
@@
and by remembering that they correspond to the cross-ratio coordinates $c_{i}$, we get the result from applying theorem 4.2. ∎
@@ 5.2 Flat cross-ratio dynamics
The flat cross-ratio dynamic was studied in [10] and [19] for constant discrete curvature $\alpha\in\mathbb{C}^{*}$, and for varying $\alpha_{i}\in\mathbb{C}^{*}$ in [16]. Hamiltonian integrability was proved in both situations.
@@
Let $n\geq 3$ and $P=((p_{1},\dots,p_{n}),M)$ be an $n$-gon in $\mathbb{P}^{1}$ such that $p_{i}\neq p_{i+1},p_{i+2}$ for each $i\in\mathbb{Z}$. Enrich it with a discrete curvature $\alpha=(\alpha_{1},\dots,\alpha_{n})$ such that $\alpha_{i}\in\mathbb{G}_{m}\setminus\{1\}$ for each $i\in\mathbb{Z}$.
@@
From [10, Cor. 2.7] we get that for an $n$-gon $P$, there are generically only two $n$-gon $Q^{+},Q^{-}$ that are $\alpha$-related to it. This gives an evolution operator in the futur/past, called the flat cross-ratio dynamics.
Next lemma is an adaptation of [10, Lemma 4.4] for non-constant $\alpha_{i}$’s.
@@
$\begin{split}c_{i}=\alpha_{i}x_{i}(1-x_{i+1}),\quad d_{i}=\alpha_{i}y_{i}(1-y_{i+1}),\\ \quad 1=x_{i}+y_{i}+x_{i}y_{i}\left(\frac{1-\alpha_{i}}{1-\alpha_{i-1}}-1\right).\end{split}$ (1)
@@
Again, it is similar to [10]. By using the “Chasles’ relation of the cross-ratio”, we have:
$c_{i}=[p_{i},p_{i+1},p_{i-1},q_{i}][p_{i},p_{i+1},q_{i},q_{i+1}][p_{i},p_{i+1},q_{i+1},p_{i+2}]=x_{i}\alpha_{i}(1-x_{i+1}),$
@@
so by referring to the permutations of the variables of the cross-ratio, we get:
@@
This rewrites as:
$1=x_{i}+y_{i}+x_{i}y_{i}\left(\frac{1-\alpha_{i}}{1-\alpha_{i-1}}-1\right).$
Conversely, the data of $x$ and $\alpha$ allows to reconstruct $y$, $c$ and $d$. We extend them to get $n$-periodic sequences. From $c$, we can choose representant $P=(p_{i})_{i\in\mathbb{Z}}$ such $c_{i}=[p_{i},p_{i+1},p_{i-1},p_{i+2}]$. Because of lemma 4.1, we know its monodromy $M_{P}$. Similarly we construct a polygon $Q=(q_{i})_{i\in\mathbb{Z}}$, with monodromy $M_{Q}$. We observe that
@@
so by a classical property of the cross-ratio and projective transformations, we necessary have $M_{P}=M_{Q}$. Because of the computation that gave us
@@
This crucial lemma allows us to define a scaling symmetry. This was already observed in [10] for constant $\alpha_{i}$, although the scaling was much simpler and the concept of scaling symmetry was not named.
@@
$\alpha_{i}(t)=1+\kappa_{i}(t\alpha_{1}-1),\quad c_{i}(t)=c_{i}\frac{\alpha_{i}(t)}{\alpha_{i}},\qquad\text{where }\kappa_{i}=\frac{1-\alpha_{i}}{1-\alpha_{1}}.$
@@
Remark 9.
If all the $\alpha_{i}$ are equal, we get back the scaling symmetry form [10, Lemma 4.4].
@@
$\kappa_{i}(t)=\frac{1-\alpha_{i}(t)}{1-\alpha_{1}(t)}=\frac{\kappa_{i}(t\alpha_{1}-1)}{(t\alpha_{1}-1)}=\kappa_{i},$
@@
Let’s now see why it commutes with the evolution operator. Take $c,d,x,y$ as in lemma 5.4. They are linked by equation 1. Note that $\frac{1-\alpha_{i}}{1-\alpha_{i-1}}$ is equal to $\frac{\kappa_{i}}{\kappa_{i-1}}$, so it stays constant when applying $\rho$. In fact, applying $\rho$ modifies $c,d,\alpha$, but leaves $x$ and $y$ unchanged. The equation 1 still holds for $c_{i}(t),d_{i}(t),\alpha_{i}(t),x_{i},y_{i}$ and so by lemma 5.4 we have $P(t)\overset{\alpha(t)}{\sim}Q(t)$. Hence $\rho$ is indeed a scaling symmetry. ∎
@@
$M^{\prime}=\sum_{i=1}^{n}\frac{(1-\alpha_{i})\alpha_{1}}{(1-\alpha_{1})\alpha_{i}}\frac{1}{p_{i}-p_{i+1}}\begin{pmatrix}p_{i}&-p_{i}p_{i+1}\\ 1&-p_{i+1}\end{pmatrix}$,
@@
$\chi_{P}(X,Y)=\sum_{i=1}^{n}\frac{(1-\alpha_{i})}{\alpha_{i}}\frac{(X-p_{i}Y)(X-p_{i+1}Y)}{p_{i}-p_{i+1}}.$
@@
$\frac{c_{i}^{\prime}(1)}{c_{i}}=\frac{(1-\alpha_{i})\alpha_{1}}{(1-\alpha_{1})\alpha_{i}},$
so by theorem 4.2 we get the formula for the infinitesimal monodromy. Still from the same theorem, we obtain that an eigenvector $(X,Y)^{T}$ is a root of the polynomial $\chi_{P}(X,Y)$. Up to multiplying it by $\frac{\alpha_{1}}{(1-\alpha_{1})}$, which we do only for cosmetic purpose, we get desired formula for the polynomial. ∎
@@ 6 A group dynamic in $\mathbb{P}^{1}$: staircase cross-ratio dynamics
We define a new polygonal dynamic in $\mathbb{P}^{1}$ given by a group of polygonal transformations. As said in the introduction, it is a generalization of other systems coming from the continuation of discrete holomorphic functions (see [12, §6] and references therein).
@@ 6.1 Definition and results
Let $\alpha,\beta\in\mathbb{G}_{m}$. Two couples of distinct points $(a,c),(b,d)$ of $\mathbb{P}^{1}$ form a $\bm{\alpha/\beta}$-quadrangle when $a,c,b,d]=\alpha/\beta$. The point $c$ is called the $\bm{\alpha/\beta}$-conjugate of $a$ with respect to $(b,d)$, and we write $c=h^{\alpha/\beta}_{b,d}(a)$. When $\alpha/\beta=-1$, the quadrangle said to be harmonic.
@@
    Projective transformations send $\alpha/\beta$-quadrangles to $\alpha/\beta$-quadrangles. Moreover, we have:
@@
    The $\alpha/\beta$-conjugation with respect to $(b,d)$ is a projective transformation $h^{\alpha/\beta}_{b,d}(z)=\frac{(\beta d-\alpha b)z-(\beta-\alpha)bd}{(\beta-\alpha)z-(\beta b-\alpha d)}$, which fixes $b,d,$ and has characteristic constant $\alpha/\beta$. Its inverse is $h^{\alpha/\beta}_{d,b}$.
@@
    If $(b,d)$ is fixed, then for $\alpha,\beta,\gamma,\delta\in\mathbb{G}_{m}$, we have $h^{\alpha/\beta}_{b,d}\circ h^{\gamma/\delta}_{b,d}=h^{\alpha\gamma/\beta\delta}_{b,d}$. Hence $h^{1}_{b,d}=id$ and $(h^{\alpha/\beta}_{b,d})^{-1}=h^{\beta/\alpha}_{b,d}$.
@@
    Let $g\in\mathbb{P}\mathrm{GL}_{2}$ which sends $(b,d)$ to $(\infty,0)$. Using the (a) and (b), we obtain
@@
We can now define the dynamic.
@@
Let $n\geq 3$ and $P=((p_{1},\dots,p_{n}),M)$ be an $n$-gon in $\mathbb{P}^{1}$ such that $p_{i}\neq p_{i+1},p_{i+2}$ for each $i\in\mathbb{Z}$. Enrich it with an $n$-periodic discrete curvature $\mu=(\mu_{i})_{i\in\mathbb{Z}}$ such that $\mu_{i}\in\mathbb{G}_{m}$ for each $i\in\mathbb{Z}$.
@@
Because of proposition 6.2(1), $\tilde{P}$ has the same monodromy as $P$. In a more compact way, we write $P\overset{\mu}{\sim}_{j}\tilde{P}$. See figure 2 for a visualisation of the action of $\phi_{j}$.
@@
Figure 2: Visualisation of the action of the flip $\phi_{j}$ at index $j$, read from up to down. It it similar to a braid on a cylinder, but it is important not to forget that it acts also on the polygon $(p_{i})_{i\in\mathbb{Z}/n\mathbb{Z}}$; indeed, the point $p_{j}$ is changed in $\tilde{p}_{j}=h^{\mu_{j}/\mu_{j-1}}_{p_{j+1},p_{j-1}}(p_{j})$.
Proposition 6.4.
@@
where the indices (and the distance) are understood modulo $n$. This is usually called the affine symmetric group (see [13]).
@@
As it was done by Adler in [3] for the polygon recutting, we will check each relations.
@@
  • $\phi_{i}\phi_{i+1}\phi_{i}=\phi_{i+1}\phi_{i}\phi_{i+1}$: This comes from the reinterpretation of a property called “3D-consistency”([30, Prop. 9]). It states that if the vertices $(p_{i-1},p_{i},p_{i+1},p_{i+2})$ and edges $(\mu_{i-1},\mu_{i},\mu_{i+1})$ are positioned on a cube like on figure 3, one can “flip around the cube” by doing $(\phi_{i+1}\phi_{i})^{3}$ and get back to the same state. So $(\phi_{i+1}\phi_{i})^{3}=\mathrm{Id}$, and since flips are involutions we get the result.∎
@@
Figure 3: A visualization of 3D consistency, meaning that one can “flip around the cube consistently”. The red vertices and edges are the ones taken originally.
@@
The group $\Gamma_{n}$ doesn’t always act faithfully. It is the case for instance when the field $k$ is finite (since any orbit is finite). However, when there is an element $q\in k$ of infinite multiplicative order [^12] For instance $2\in\mathbb{Q}$., then each iterate of $\psi:=\phi_{n}\circ\dots\circ\phi_{1}\in\Gamma_{n}$ acts differently (see subsection 6.2). Hence the dynamic doesn’t consist only of finites orbits, which is coherent with our conjectures.
@@
$c_{i}=\tilde{c}_{i}\quad\forall i\neq j-2,j-1,j,j+1,$ ($\star_{i}$)
@@
$c_{i}=\tilde{c}_{i}\quad\forall i\neq j-2,j-1,j,j+1.$
@@
Conversely, choose some representants $P=(p_{i})_{i\in\mathbb{Z}}$ and $\tilde{P}=(\tilde{p}_{i})_{i\in\mathbb{Z}}$, equipped with discrete curvatures $\mu$ and $\tilde{\mu}$, such that $c_{i}=[p_{i},p_{i+1},p_{i-1},p_{i+2}]$ and $\tilde{c}_{i}=[\tilde{p}_{i},\tilde{p}_{i+1},\tilde{p}_{i-1},\tilde{p}_{i+2}]$. Because of lemma 4.1, we know their monodromies $M_{P}$, $M_{\tilde{P}}$.
@@
By doing almost the same and using $\star_{j-2}$ instead, we obtain that:
@@
This gives us $M_{P}=M_{\tilde{P}}$, because they both map $(p_{j-n+1},p_{j-n+2},p_{j-n+3})$ to $(p_{j+1},p_{j+2},p_{j+3})$. Hence we have that:
@@
By rearranging the equation $\star_{j}$ in the same way as in the beginning of this proof, we see that indeed:
@@
Considering the scaling symmetries for each $\eta\in\mathbb{G}_{m}$ is useful. Indeed, if we consider the discrete curvature $\lambda\mu$ for $\lambda\in\mathbb{G}_{m}$, then we have the following commutative diagram.
$(c,\mu)$$(c^{(\eta)}(t),\mu^{(\eta)}(t))$$(c,\lambda\mu)$$(c^{(\lambda\eta)}(t),\lambda\mu^{(\lambda\eta)}(t))$$\scriptstyle{\rho^{(\eta)}(t)}$$\scriptstyle{\lambda}$$\scriptstyle{\lambda}$$\scriptstyle{\rho^{(\lambda\eta)}(t)}$
However, we’ll see later that the eigenlines of the infinitesimal monodromy don’t depend on the choice of $\eta$.
@@
Then, because of the converse part of lemma 6.5, we need to check that equations $\star_{i}$, $\star_{j-2}$, $\star_{j-1}$, $\star_{j}$ and $\star_{j+1}$ hold for the pairs $(c^{(\eta)}(t),\mu^{(\eta)}(t))$ and $(\tilde{c}^{(\eta)}(t),\tilde{\mu}^{(\eta)}(t))$.

### v3

@@ 5 Two examples of dynamics in $\mathbb{P}^{1}$
We apply the results of the previous section to three kinds of polygonal dynamics: the leapfrog map, the “flat” cross-ratio and the newly defined “stair-case” cross-ratio dynamics. The two last ones are linked to the theory of discrete holomorphic functions (see [affolterSchwarzianOctahedronRecurrence2024, §6]) and are endowed with a discrete curvature, which is an $n$-periodic sequence $\mu=(\mu_{i})_{i\in\mathbb{Z}}$, where each $\mu_{i}\in\mathbb{G}_{m}$ should be thought to live on the edge linking $p_{i}$ and $p_{i+1}$. They are left unchanged by the action of $\mathbb{P}\mathrm{GL}_{2}$, but can be modified by the scaling symmetry.
@@ 5.1 The leapfrog map
The leapfrog map is defined in [gekhtmanIntegrableClusterDynamics2016, §5.2] as an analogue of the pentagram map in $\mathbb{P}^{1}$, and the authors notice that it seems to be the hyperbolic counterpart of the polygon recutting defined in [adlerRecuttingsPolygons1993] (which would be the euclidean case). For a matter of internal coherence, the notations of [gekhtmanIntegrableClusterDynamics2016] are changed to fit the ones from here.
@@
with the same monodromy, which we intertwine this way
@@
There is a scaling symmetry acting as
@@
We express the system of coordinates $(o^{-},o)$ with the help of two other systems, and then get the result. According to [gekhtmanIntegrableClusterDynamics2016, Prop. 5.12], we have the coordinate systems $(x,y)$ and $(p,r)$, defined by:
@@
They are linked by the relations [^14] The original paper forgot the minus signs, but it doesn’t impact their reasoning.
@@
But according to [gekhtmanIntegrableClusterDynamics2016, Rmk. 3.5], there is a scaling symmetry (actually working for all of their generalised pentagram maps) defined for any $t\in\mathbb{G}_{m}$ by
@@
$M^{\prime}(1)=\sum_{i=1}^{n}\frac{1}{s_{i}^{-}-s_{i}}\begin{pmatrix}s_{i}^{-}&-s_{i}^{-}s_{i}\\ 1&-s_{i}\end{pmatrix}$,
@@
and by remembering that they correspond to the cross-ratio coordinates $c_{i}$, we get the result by applying theorem 4.2. ∎
@@ 5.2 Flat cross-ratio dynamics
The flat cross-ratio dynamics arise from the continuation of discrete holomorphic functions, for which the geometry of the initial condition (meaning $P$) is flat. See figure 2 and compare with the staircase cross-ratio dynamics depicted on figure 3.
It was studied in [arnoldCrossratioDynamicsIdeal2022] and [hetrich-jerominPeriodicDiscreteConformal2001] for constant discrete curvature $\alpha\in\mathbb{C}\setminus\{0,1\}$, where they respectively prove Hamiltonian and algebraic integrability. It was generalized for varying $(\alpha_{i})\in(\mathbb{C}\setminus\{0,1\})^{n}$ in [affolterIntegrableDynamicsProjective2023].
@@
Let $n\geq 3$ and $P=((p_{1},\dots,p_{n}),M)$ be an $n$-gon in $\mathbb{P}^{1}$. Enrich it with a discrete curvature $\alpha=(\alpha_{1},\dots,\alpha_{n})$ such that $\alpha_{i}\in\mathbb{G}_{m}\setminus\{1\}$ for each $i\in\mathbb{Z}$.
@@
From [arnoldCrossratioDynamicsIdeal2022, Cor. 2.7] we get that for an $n$-gon $P$, there are generically only two $n$-gon $Q^{+},Q^{-}$ that are $\alpha$-related to it. This gives an evolution operator in the futur/past, called the flat cross-ratio dynamics.
[…]
Figure 2: Flat cross-ratio dynamics seen in the context of discrete holomorphic functions. The 66-gon $P=((p_{1},\dots,p_{6}),M)$ is $\alpha$-related in the past to $Q^{-}$ and in the future to $Q^{+}$.
Next lemma is an adaptation of [arnoldCrossratioDynamicsIdeal2022, Lemma 4.4] for non-constant $\alpha_{i}$’s.
@@
$\begin{split}c_{i}=\alpha_{i}x_{i}(1-x_{i+1}),\quad d_{i}=\alpha_{i}y_{i}(1-y_{i+1}),\\ \quad 1=x_{i}+y_{i}+x_{i}y_{i}\left(\frac{\alpha_{i}-1}{\alpha_{i-1}-1}-1\right).\end{split}$ (3)
@@
Again, it is similar to [arnoldCrossratioDynamicsIdeal2022]. By using “Chasles’ relation of the cross-ratio”, we have
$c_{i}$ $=[p_{i},p_{i+1},p_{i-1},q_{i}][p_{i},p_{i+1},q_{i},q_{i+1}][p_{i},p_{i+1},q_{i+1},p_{i+2}]$
$=x_{i}\alpha_{i}(1-x_{i+1}),$
@@
so by referring to the permutations of the variables of the cross-ratio, we get
@@
This rewrites as
$1=x_{i}+y_{i}+x_{i}y_{i}\left(\frac{\alpha_{i}-1}{\alpha_{i-1}-1}-1\right).$
Conversely, the data of $x$ and $\alpha$ allows to reconstruct $y$, $c$ and $d$. We extend them to get $n$-periodic sequences. From $c$, we can choose representative $P=(p_{i})_{i\in\mathbb{Z}}$ such $c_{i}=[p_{i},p_{i+1},p_{i-1},p_{i+2}]$. Because of lemma 2.6, we know its monodromy $M_{P}$. Similarly we construct a polygon $Q=(q_{i})_{i\in\mathbb{Z}}$, with monodromy $M_{Q}$. We observe that
@@
hence we have $M_{P}=M_{Q}$. Because of the computation that gave us
@@
This crucial lemma allows us to define a scaling symmetry. This was already observed in [arnoldCrossratioDynamicsIdeal2022, Lemma 4.4] for constant $\alpha$, although the scaling was much simpler and the concept of scaling symmetry was not named.
@@
$\alpha_{i}(t)=1+\kappa_{i}(t\alpha_{1}-1),\quad c_{i}(t)=c_{i}\frac{\alpha_{i}(t)}{\alpha_{i}},\qquad\text{where }\kappa_{i}=\frac{\alpha_{i}-1}{\alpha_{1}-1}.$
@@
$\kappa_{i}(t)=\frac{\alpha_{i}(t)-1}{\alpha_{1}(t)-1}=\frac{\kappa_{i}(t\alpha_{1}-1)}{(t\alpha_{1}-1)}=\kappa_{i},$
@@
Let’s now see why it commutes with the evolution operator. Take $c,d,x,y$ as in lemma 5.4. They are linked by equation 3. Note that $\frac{\alpha_{i}-1}{\alpha_{i-1}-1}$ is equal to $\frac{\kappa_{i}}{\kappa_{i-1}}$, so it stays constant when applying $\rho$. In fact, applying $\rho$ modifies $c,d,\alpha$, but leaves $x$ and $y$ unchanged. Equation 3 still holds for $c_{i}(t),d_{i}(t),\alpha_{i}(t),x_{i},y_{i}$ and so by lemma 5.4 we have $P(t)\overset{\alpha(t)}{\sim}Q(t)$. Hence $\rho$ is indeed a scaling symmetry. ∎
@@
$M^{\prime}(1)=\sum_{i=1}^{n}\frac{(\alpha_{i}-1)\alpha_{1}}{(\alpha_{1}-1)\alpha_{i}}\frac{1}{p_{i}-p_{i+1}}\begin{pmatrix}p_{i}&-p_{i}p_{i+1}\\ 1&-p_{i+1}\end{pmatrix}$,
@@
$\chi_{P}(X,Y)=\sum_{i=1}^{n}\frac{(\alpha_{i}-1)}{\alpha_{i}}\frac{(X-p_{i}Y)(X-p_{i+1}Y)}{p_{i}-p_{i+1}}.$
@@
$\frac{c_{i}^{\prime}(1)}{c_{i}(1)}=\frac{(\alpha_{i}-1)\alpha_{1}}{(\alpha_{1}-1)\alpha_{i}},$
so by theorem 4.2 we get the formula for the infinitesimal monodromy. Still from the same theorem, we obtain that an eigenvector $(X,Y)^{T}$ is a root of the polynomial $\chi_{P}(X,Y)$. Up to multiplying it by $\frac{\alpha_{1}-1}{\alpha_{1}}$, which we do only for cosmetic purpose, we get desired formula for the polynomial. ∎
@@ 6 A group dynamic in $\mathbb{P}^{1}$: staircase cross-ratio dynamics
[…]
Figure 3: Staircase cross-ratio dynamics seen in the context of discrete holomorphic functions. Applying the flip $\phi_{3}$ on the 66-gon $P=((p_{1},\dots,p_{6}),M)$ changes $p_{3}$ to $\tilde{p}_{3}$ and interchanges $\mu_{2}$ with $\mu_{3}$.
We define a new polygonal dynamic in $\mathbb{P}^{1}$ given by a group of polygonal transformations called flips. [^15] It most certainly admits an interpretation as cluster transformations. The staircase cross-ratio dynamics arise from the continuation of discrete holomorphic functions, for which the geometry of the initial condition (meaning $P$) is staircase-like. See figure 3 and compare with the flat cross-ratio dynamics depicted on figure 2.
However, this dynamic is more general than the problem of continuation of discrete holomorphic functions, for which some flips are not allowed by the geometry. [^16] On the example presented on figure 3, the flip $\phi_{5}$ is not allowed since $p_{5}$ is not on a “corner” like $p_{3}$. The more abstract way to consider the dynamic is to forget the underlying geometry and to see the group as a braid group, as on figure 4 and theorem 6.4.
@@ 6.1 Definition and results
Let $\alpha,\beta\in\mathbb{G}_{m}$. Two couples of distinct points $(a,c),(b,d)$ of $\mathbb{P}^{1}$ form a $\bm{\alpha/\beta}$-quadrangle when $a,c,b,d]=\alpha/\beta$. The point $c$ is called the $\bm{\alpha/\beta}$-conjugate of $a$ with respect to $(b,d)$, and we write
$c=h^{\alpha/\beta}_{b,d}(a).$
When $\alpha/\beta=-1$, the quadrangle said to be harmonic.
@@
    Projective transformations send $\alpha/\beta$-quadrangles to $\alpha/\beta$-quadrangles. Moreover, we have
@@
    The $\alpha/\beta$-conjugation with respect to $(b,d)$ is a projective transformation
    $h^{\alpha/\beta}_{b,d}(z)=\frac{(\beta d-\alpha b)z-(\beta-\alpha)bd}{(\beta-\alpha)z-(\beta b-\alpha d)},$
    which fixes $b$ and $d$, and has characteristic constant $\alpha/\beta$. Its inverse is $h^{\alpha/\beta}_{d,b}$.
@@
    If $(b,d)$ is fixed, then for $\alpha,\beta,\gamma,\delta\in\mathbb{G}_{m}$ we have
    $h^{\alpha/\beta}_{b,d}\circ h^{\gamma/\delta}_{b,d}=h^{\alpha\gamma/\beta\delta}_{b,d}.$
    Hence $h^{1}_{b,d}=\mathrm{Id}$ and $(h^{\alpha/\beta}_{b,d})^{-1}=h^{\beta/\alpha}_{b,d}$.
@@
    Let $g\in\mathbb{P}\mathrm{GL}_{2}$ which sends $(b,d)$ to $(\infty,0)$. Using (a) and (b), we obtain
@@
We can now define the dynamic. See figure 4 for a visualisation of it.
@@
Let $n\geq 3$ and $P=((p_{1},\dots,p_{n}),M)$ be an $n$-gon in $\mathbb{P}^{1}$. Enrich it with an $n$-periodic discrete curvature $\mu=(\mu_{i})_{i\in\mathbb{Z}}$ such that $\mu_{i}\in\mathbb{G}_{m}$ for each $i\in\mathbb{Z}$.
@@
We write $P\overset{\mu}{\sim}_{j}\tilde{P}$.
@@
Figure 4: Visualisation of the action of the flip $\phi_{j}$ at index $j$, read from down to up. It it similar to a braid on a cylinder, but it is important to keep in mind that it also acts on the vertices; indeed, the point $p_{j}$ is changed in $\tilde{p}_{j}=h^{\mu_{j}/\mu_{j-1}}_{p_{j+1},p_{j-1}}(p_{j})$.
Theorem 6.4.
@@
where the indices (and the distance) are understood modulo $n$. This is usually called the affine symmetric group (see [lewisAffineSymmetricGroup2021]).
@@
As it was done by Adler in [adlerRecuttingsPolygons1993] for the polygon recutting, we will check each relations.
@@
  • $\phi_{i}\phi_{i+1}\phi_{i}=\phi_{i+1}\phi_{i}\phi_{i+1}$: This comes from the reinterpretation of a property called “3D-consistency” ([bobenko_discrete_2008, Prop. 9]). It states that if the vertices $(p_{i-1},p_{i},p_{i+1},p_{i+2})$ and edges $(\mu_{i-1},\mu_{i},\mu_{i+1})$ are positioned on a cube like on figure 5, one can “flip around the cube” by doing $(\phi_{i+1}\phi_{i})^{3}$ and get back to the same state. So $(\phi_{i+1}\phi_{i})^{3}=\mathrm{Id}$, and since flips are involutions we get the result.∎
@@
Figure 5: A visualization of 3D consistency. This means that starting from the red vertices and edges, doing the sequences of flips $\phi_{i}\phi_{i+1}\phi_{i}$ or $\phi_{i+1}\phi_{i}\phi_{i+1}$ give the same result.
@@
The group $\Gamma_{n}$ doesn’t always act faithfully, for instance when the field $F$ is finite (since any orbit is finite). However, when there is an element $q\in F$ of infinite multiplicative order [^17] For instance $2\in\mathbb{Q}$., then each iterate of $\psi:=\phi_{n}\circ\dots\circ\phi_{1}\in\Gamma_{n}$ acts differently (see subsection 6.2). Hence the dynamic doesn’t consist only of finite orbits.
@@
$c_{i}$ $=\tilde{c}_{i}\qquad\quad\forall i\neq j-2,j-1,j,j+1,$ ($\star_{i}$)
@@
$c_{i}=\tilde{c}_{i}\qquad\forall i\neq j-2,j-1,j,j+1.$
@@
Conversely, choose some representatives $P=(p_{i})_{i\in\mathbb{Z}}$ and $\tilde{P}=(\tilde{p}_{i})_{i\in\mathbb{Z}}$, equipped with discrete curvatures $\mu$ and $\tilde{\mu}$, such that $c_{i}=[p_{i},p_{i+1},p_{i-1},p_{i+2}]$ and $\tilde{c}_{i}=[\tilde{p}_{i},\tilde{p}_{i+1},\tilde{p}_{i-1},\tilde{p}_{i+2}]$. Because of lemma 2.6, we know their monodromies $M_{P}$, $M_{\tilde{P}}$.
@@
By doing almost the same and using $\star_{j-2}$ instead, we obtain
@@
This gives us $M_{P}=M_{\tilde{P}}$, because they both map $(p_{j-n+1},p_{j-n+2},p_{j-n+3})$ to $(p_{j+1},p_{j+2},p_{j+3})$. Hence we have
@@
By rearranging the equation $\star_{j}$ in the same way as in the beginning of this proof, we see that
@@
Considering the scaling symmetries for each $\eta\in\mathbb{G}_{m}$ is useful. Indeed, for any $\lambda\in\mathbb{G}_{m}$, the discrete curvatures $\mu$ and $\lambda\mu$ provide the same dynamic. They are related by this commutative diagram.
$(c,\mu)$$\rho^{(\eta)}(t)\cdot(c,\mu)$$(c,\lambda\mu)$$\rho^{(\eta/\lambda)}(t)\cdot(c,\lambda\mu)$$\scriptstyle{\rho^{(\eta)}(t)}$$\scriptstyle{\lambda}$$\scriptstyle{\lambda}$$\scriptstyle{\rho^{(\eta/\lambda)}(t)}$
We will see later that the infinitesimal monodromy is independent of the choice of $\eta$.
@@
Let’s show that $y_{j}^{(\eta)}(t)$ and $x_{j}^{(\eta)}(t)$ are constant. Since both cases are similar, we only perform the computation for $y_{j}$. To do so, take the lift [^18] See remark 5 for details. such that for any $t$
$(p_{j-1}^{(\eta)}(t),p_{j}^{(\eta)}(t),p_{j+1}^{(\eta)}(t))=(1,\infty,0),$
which forces that $p_{j+2}^{(\eta)}(t)=c_{j}^{(\eta)}(t)$. By proposition 6.2(2), we have
$\tilde{p}_{j}^{(\eta)}(t)=\frac{\mu_{j-1}^{(\eta)}(t)}{\mu_{j-1}^{(\eta)}(t)-\mu_{j}^{(\eta)}(t)}.$
From this we obtain
$y_{j}^{(\eta)}(t)$ $=[0,p_{j+2}^{(\eta)}(t),\infty,\tilde{p}_{j}^{(\eta)}(t)]$
$=\frac{\tilde{p}_{j}^{(\eta)}(t)-p_{j+2}^{(\eta)}(t)}{\tilde{p}_{j}^{(\eta)}(t)}$
$=1-c_{j}\frac{t\mu_{j}}{\mu_{j}^{(\eta)}(t)}\frac{\mu_{j-1}^{(\eta)}(t)-\mu_{j}^{(\eta)}(t)}{\mu_{j-1}^{(\eta)}(t)}$
$=1-c_{j}\mu_{j}(\frac{1-(t-1)\eta\mu_{j}}{\mu_{j}}-\frac{1-(t-1)\eta\mu_{j-1}}{\mu_{j-1}})$
$=1-c_{j}\mu_{j}(\frac{1}{\mu_{j}}-\frac{1}{\mu_{j-1}}),$
so it is constant. Then, because of the converse part of lemma 6.5, we need to check that equations $\star_{i}$, $\star_{j-2}$, $\star_{j-1}$, $\star_{j}$ and $\star_{j+1}$ hold for the pairs $(c^{(\eta)}(t),\mu^{(\eta)}(t))$ and $(\tilde{c}^{(\eta)}(t),\tilde{\mu}^{(\eta)}(t))$.